# Kevin Lin3, Xiaofei Wang3, Zhengyuan Yang3, Linjie Li3, Chung-Ching Lin3, Jianwen Xie4

### v1

@@ Abstract
To address this, we introduce EdiVal-Agent , an automated, scalable, and fine-grained evaluation framework for multi-turn instruction-based editing from an object-centric perspective, supported by a suite of expert tools. Given an image, EdiVal-Agent  first decomposes it into semantically meaningful objects, then synthesizes diverse, context-aware editing instructions. For evaluation, it integrates VLMs with open-vocabulary object detectors to assess instruction following, uses semantic-level feature extractors to evaluate content consistency, and leverages human preference models to judge visual quality. We show that combining VLMs with object detectors yields stronger agreement with human judgments in instruction-following evaluation compared to using VLMs alone and CLIP-based metrics. Furthermore, the pipeline’s modular design allows future tools to be seamlessly integrated, enhancing evaluation accuracy over time.
Instantiating this pipeline, we build EdiVal-Bench, a multi-turn editing benchmark covering 9 instruction types and 11 state-of-the-art editing models spanning autoregressive (AR), flow-matching, and diffusion paradigms. Our analysis shows that Nano Banana achieves the best overall balance of instruction following and content consistency while maintaining surprisingly low latency; GPT-Image-1 exhibits the strongest instruction following across turns but ranks among the lowest in content consistency and has high latency; AR models substantially outperform non-AR models in multi-turn editing by better preserving contextual coherence; and Qwen-Image-Edit is strong on the first turn but degrades in later turns due to exposure bias. We demonstrate that EdiVal-Agent can be used to identify existing failure modes, thereby informing the development of the next generation of editing models. Project page: https://tianyucodings.github.io/EdiVAL-page/
Figure 1: Overview of our workflow and representative model’s performance. For visualization, we adopt two thresholds: a consistency score of at least $90$ and a visual quality score of at least 66. Details of the automated evaluation pipeline are provided in Figure 2 and Section 3. In multi-turn editing, models exhibit distinct weaknesses: GPT-Image-1 struggles with content consistency, Qwen-Image-Edit underperforms in both visual quality and content consistency, and FLUX.1-Kontext-dev lags in instruction following, whereas Nano Banana shows no single dominant weakness. A comprehensive analysis is presented in Sec. 5 and Tab. 2.
@@ 1 Introduction
What truly defines the success of an instruction-based image editor? At its core, editing requires making targeted, instruction-driven changes while preserving contextual consistency and perceptual realism—often across multiple refinement turns. Yet current evaluation practice struggles to capture this multifaceted objective.
@@
When ground-truth edited images are available, a common strategy is to compare model outputs against these references Zhang et al. (2023); Zhao et al. (2024); Yu et al. (2025); Sheynin et al. (2024). Typical metrics include pixel-level distances (e.g., L1/L2) and semantic similarities (e.g., DINO Caron et al. (2021) and CLIP Radford et al. (2021)). While informative, such metrics suffer from two structural issues: (i) the space of acceptable edits is inherently large, whereas a single reference provides only one realization; and (ii) references are frequently synthesized by existing editing models (e.g., Prompt-to-Prompt Hertz et al. (2023), SDXL Podell et al. (2024), DALLE-2 Ramesh et al. (2022)), thereby importing their biases and limitations into the evaluation itself. Consequently, high reference similarity does not necessarily imply faithful instruction following, preservation of irrelevant content, or aesthetically plausible outcomes.
@@
A complementary line of work employs VLMs as interpretable evaluators—for example, VIEScore Ku et al. (2023), HQ-Edit Hui et al. (2024), and Complex-Edit Yang et al. (2025)—and queries VLMs about specific aspects of an edit. While VLMs offer holistic, language-mediated judgments, they remain insufficient for precise editing assessment for several reasons. First, for instruction-following evaluation, they are notoriously poor at spatial reasoning Zhang et al. (2025b); Chen et al. (2024); Chang et al. (2025) and are prone to object hallucinations in existence, category, attributes, and relations Bai et al. (2024). These issues together undermine their ability to assess common object-related edit instructions. Second, they have limited sensitivity to pixel-level changes and frequently miss subtle, localized modifications Vo et al. (2025) (e.g., fine structures, small attribute shifts, etc.), which are crucial for evaluating content consistency. Third, since they are predominantly pretrained on natural images rather than synthetic generations, their priors are miscalibrated for artifacts and aesthetics, leading to failures in detecting common generative defects (e.g., extra fingers) and in modeling perceptual “naturalness” Liang et al. (2024); Xu et al. (2023); Ma et al. (2025a), which humans are sensitive to. Consequently, VLM-only scoring lacks the precision and reliability required for fine-grained evaluation across instruction following, content consistency, and visual quality.
@@
Figure 2: Framework of EdiVal-Agent. It first decomposes images into semantically meaningful objects, such as metal yellow sign and metal brown pole, and identifies their contextual relationships, e.g., they are both in foreground. It then generates diverse and proper editing scenarios at scale which are based on the initial analysis, e.g., Change the color of metal brown pole to gray. Finally, it systematically evaluates editing model outputs from multiple axes, including instruction following, content consistency, and visual quality, specifically by integrating VLMs (Qwen2.5-VL) with open-vocabulary object detectors (Grouding-DINO) to assess instruction following, using semantic-level feature extractors (DINOv3) and pixel-level metrics ($L_{1}$ distance) to evaluate content consistency, and leveraging human preference models (HPSv3) to judge visual quality. Our agentic pipeline is agnostic to the expert tools used and can be readily enhanced with more advanced tools in the future.
@@
To address these challenges, we introduce EdiVal-Agent: an automated, scalable, and fine-grained evaluation agent for multi-turn instruction-based image editing from an object-centric perspective, supported by a suite of expert tools. As shown in Fig. 2, EdiVal-Agent  first decomposes images into semantically meaningful components and identifies their contextual relationships. It then generates diverse and proper editing scenarios at scale, which are based on the initial analysis. Finally, it systematically evaluates editing model outputs from multiple axes, specifically by integrating VLMs with open-vocabulary object detectors to assess instruction following, using semantic-level feature extractors to evaluate content consistency, and leveraging human preference models to judge visual quality. We show that combining VLMs with object detectors yields stronger agreement with human judgments in instruction-following evaluation compared to thresholded CLIP directional (CLIP_dir) scores Gal et al. (2022) and using VLMs alone, as evidenced in Sec. 4. Note that the full evaluation process, from evaluation instruction creation to evaluation metrics, is fully automated, without human intervention. This results in a scalable pipeline that can be easily expanded.
@@
Instantiating the agentic pipeline, we curate a new multi-turn image editing benchmark, EdiVal-Bench, featuring 9 instruction types and 11 existing editing models—spanning autoregressive (AR), flow-matching, and diffusion paradigms, across both closed- and open-source systems—conduct fine-grained analyses, and draw actionable insights. Empirically, as demonstrated in Fig. 1 and Tab. 2, GPT-Image-1 excels at instruction following yet ranks near the bottom in content consistency, whereas Nano Banana performs strongly on both axes. Besides, recently released open-sourced models like Qwen-Image-Edit significantly degrade in visual quality when editing turns increase, and fails to maintain consistency with the base image, while FLUX.1-Kontext-dev lags in instruction following. We further study editing behavior differences between AR and non-AR models in the multi-turn setting, and contrast multi-turn editing with complex single-shot prompts Yang et al. (2025), highlighting complementary strengths and failure modes. We hope that our agent pipeline, benchmark, and analyses accelerate the transition of multi-turn editing toward practical applications.
@@
Key contributions. 1) Agent: EdiVal-Agent is a fully automated, scalable evaluator that performs object-centric decomposition, generates multi-turn editing instructions, and integrates VLM reasoning with symbolic pixel- and semantic-level metrics via grounding tools; 2) Benchmark: using EdiVal-Agent, we build EdiVal-Bench with 1,716 instructions across 9 types and 3 turns on 572 real-world images; 3) Human agreement: EdiVal-Agent achieves 81.3% agreement with human ratings in instruction following, outperforming zero-shot VLMs and CLIP-based methods; 4) Evaluation: we assess 11 editors (diffusion, flow-matching, autoregressive) along instruction following, content consistency, and visual quality; 5) Insights: AR models better preserve contextual coherence and instruction accuracy across turns, while non-AR models exhibit exposure bias and cumulative degradation; 6) Artifacts & settings: we reveal luminance drift across turns, and contrast multi-turn against complex single-shot editing to delineate strengths and weaknesses across model families.
2 Related Work
Instruction-based editing models.
InstructPix2Pix (IP2P) Brooks et al. (2023) introduced a two-stage recipe that converts a text-to-image diffusion model Rombach et al. (2022); Zhang et al. (2025a) into an editor: (i) synthesize paired editing data using Stable Diffusion Rombach et al. (2022) and training-free techniques such as Prompt-to-Prompt Hertz et al. (2023); (ii) fine-tune the diffusion model on these pairs. Subsequent systems—MagicBrush Zhang et al. (2023), UltraEdit Zhao et al. (2024), and AnyEdit Yu et al. (2025)—scale this paradigm to large, fine-grained real-image editing. More recent work (e.g., OmniGen Xiao et al. (2025); Wu et al. (2025b), Step1X-Edit Liu et al. (2025), FLUX.1 Kontext Labs et al. (2025), and Qwen-Image-Edit Wu et al. (2025a)) adopts task-aware architectures and increasingly leverages flow matching Liu et al. (2022); Lipman et al. (2022); Zhang et al. (2024).
Editing evaluation.
Recent work relies exclusively on VLMs as interpretable judges—e.g., VIEScore Ku et al. (2023), HQ-Edit Hui et al. (2024), and Complex-Edit Yang et al. (2025)—by querying models such as GPT-4o OpenAI (2025) about specific aspects of an edit. While VLMs offer holistic, language-mediated assessments, they are insufficient on their own: they are notoriously poor at spatial reasoning Zhang et al. (2025b); Cheng et al. (2024); Chen et al. (2024); Qharabagh et al. (2024); Chang et al. (2025) and are prone to object hallucinations in existence, category, attributes, and relations Bai et al. (2024); they have limited sensitivity to pixel-level changes and frequently miss subtle, localized modifications Vo et al. (2025) (e.g., fine structures, small attribute shifts, etc.), which are crucial for evaluating content consistency; they are miscalibrated for artifacts and aesthetics Liang et al. (2024); Xu et al. (2023); Ma et al. (2025a), which humans are sensitive to. Our approach, EdiVal-Agent, addresses these gaps by integrating VLM-based reasoning with grounding tools, symbolic, object-centric pixel- and semantic-level tools, and human preference models, yielding a precise and interpretable evaluation of instruction-based editing.
Editing tasks.
[…]
@@ 3.1 Overview
EdiVal-Agent is a fully automated, scalable framework for evaluating instruction-based image editing from a symbolic, object-centric perspective, as illustrated in Fig. 2. The pipeline comprises three stages: (1) Decomposition uses a VLM (e.g., GPT-4o; other VLMs are viable alternatives) to extract structured, object-level descriptions—objects, attributes, and relations—enabling symbolic reasoning; (2) Instruction Generation produces multi-turn, diverse, compositional prompts by maintaining an explicit object pool and sampling from nine instruction types spanning subject-, attribute-, relational-, text-, count-, and global-level edits; (3) Evaluation scores edited images along Instruction Following, Content Consistency, and Visual Quality, combining Grounding-DINO Liu et al. (2024a), Qwen-VL 2.5 Yang et al. (2024), DINOv3 Siméoni et al. (2025), pixel-wise $L_{1}$, and the HPSv3 quality scorer Ma et al. (2025a). Our agentic pipeline is agnostic to the agent VLM/expert tools used and can be readily enhanced with more advanced tools in the future.
@@ 3.4 Step 3: Evaluation
Instruction Following separates types by evaluability. Symbolically verifiable types—subject_add, subject_remove, subject_replace, position_change, count_change—are solely checked with Grounding-DINO Liu et al. (2024b) using prompts derived from the instruction and object descriptors; compliance uses detection confidence, counts, and geometry (existence and relative positions), following GENEVAL Ghosh et al. (2023) and T2I-CompBench Huang et al. (2023). Semantically verifiable types—color_alter, material_alter, text_change, and background_change—are evaluated with Qwen2.5-VL Yang et al. (2024) using instruction-specific templates (See details in Appendix), applied to detector-guided (i.e., Grounding DINO) object crops to focus on the edited regions. The multiturn instruction following accuracy aggregates per-instruction pass/fail along the chain (see details in Appendix).
Content Consistency assesses whether non-target content is preserved at semantic and pixel levels. Let $\mathcal{B}^{t}$ denote the Grounding-DINO detections at turn $t$; define object and background regions as $\Omega^{t}_{\text{obj}}$ and $\Omega^{t}_{\text{bg}}$. For objects, we crop each element of $\mathcal{P}^{\text{unch}}_{t}$ from $I^{0}$ and $I^{t}$ and compute DINOv3 cosine similarity $s^{\text{obj}}=\cos\!\big{(}\phi(I^{0}_{\text{obj}}),\phi(I^{t}_{\text{obj}})\big{)}$ and a normalized pixel metric $q^{\text{obj}}=1-\|I^{0}_{\text{obj}}-I^{t}_{\text{obj}}\|_{1}/Z$, averaged over unchanged objects (larger is better). Here $\|\cdot\|_{1}$ sums over all pixels and channels within the crop, and $Z$ normalizes the pixelwise L1 distance to $[0,1]$; for a region with $C$ channels, $|\Omega|$ pixels, and per-channel dynamic range $\Delta$ (e.g., $\Delta=255$ for 8-bit RGB or $\Delta=1$ for $[0,1]$-scaled floats), we set $Z=C\,|\Omega|\,\Delta$. For background, we mask $\mathcal{B}^{t}$ and compute the same two measures over $\Omega^{0}_{\text{bg}}$ versus $\Omega^{t}_{\text{bg}}$; if a background_change occurs at turn $t$, background scoring is disabled thereafter and omitted from the combined score.
Visual Quality is measured using Human Preference Score v3 (HPSv3) Ma et al. (2025b), which yields scores aligned with human preference and captures both aesthetics and visual plausibility for image synthesis. We avoid general-purpose VLMs for aesthetic judgments because they are biased toward real-image semantics and less sensitive to synthesis artifacts. We also observe systematic overexposure in some models under multiturn editing; to quantify this effect, we also compute low-level exposure statistics (e.g., luminance percentiles).
@@
We omit style_change from the current taxonomy. Open-vocabulary detectors optimized for natural images (e.g., Grounding-DINO) generalize poorly to strongly stylized domains, undermining reliable symbolic verification. Extending EdiVal-Agent with style-aware recognition is promising future work. After language-based extraction, we validate objects with Grounding-DINO, pruning low-confidence or ambiguous detections, and canonicalizing names to a uniform schema. This stabilizes the object pools, reduces error propagation during instruction generation, and tightens the coupling between symbolic references and detector-visible entities. The pipeline is fully automated and scales linearly with dataset size; privacy is enforced by excluding sensitive images at ingestion and never recording identities—only attributes of visible clothing and accessories.
4 Measuring Human Agreement
Figure 3: Results of human agreement. Dashed lines represent the average accuracy of each method. EdiVal-Agent achieves 81.3% human agreement accuracy, significantly outperforming the VLM (Qwen2.5-VL) at 75.2% and thresholded CLIP_dir at 65.4%. Note that the CLIP_dir threshold is tuned separately for each task. The inner-annotation baselines indicate the ceiling performance achievable by the tools.

### v2

# EdiVal-Agent: An Object-Centric Framework for Automated, Fine-Grained Evaluation of Multi-Turn Editing

@@ Abstract
To address this, we introduce EdiVal-Agent, an automated and fine-grained evaluation framework grounded in an object-centric perspective, designed to assess not only standard single-turn but also multi-turn instruction-based editing with precision. Given an input image, EdiVal-Agent first decomposes it into semantically meaningful objects, then synthesizes diverse, context-aware editing instructions while dynamically updating object pools across turns. These two stages enable two novel object-centric metrics tailored for multi-turn evaluation and one global metric of visual quality: 1) EdiVal-IF, which measures instruction following by combining open-vocabulary object detectors for symbolic checks with VLMs for semantic verification on detector-guided crops; 2) EdiVal-CC, which evaluates content consistency by calculating semantic similarity of unchanged objects and background using the evolving object pools; and 3) EdiVal-VQ, which quantifies changes in overall visual quality with human preference models.
Instantiating this pipeline, we build EdiVal-Bench, a multi-turn editing benchmark covering 9 instruction types and 13 state-of-the-art editing models spanning in-context  [^1] In this paper, we label certain closed-source models—GPT-Image-1, Nano Banana, and Gemini 2.0 Flash Image—as in-context, since they are integrated into autoregressive language models in the web UI and support in-context multi-turn editing., flow-matching, and diffusion paradigms. We further conduct experiments comparing multi-turn editing with single-shot complex editing, highlighting the distinctive characteristics of different model paradigms. We demonstrate that EdiVal-Agent can be used to identify existing failure modes, thereby informing the development of the next generation of editing models. Project page: https://tianyucodings.github.io/EdiVAL-page/
Figure 1: Overview of our workflow and representative model’s performance. For visualization, we adopt two thresholds: a consistency score of at least $90$ and a visual quality score of at least 66. Details of the automated evaluation pipeline are provided in Figure 2 and Section 2. In multi-turn editing, models exhibit distinct weaknesses: GPT-Image-1 struggles with content consistency, Qwen-Image-Edit underperforms in both visual quality and content consistency, and FLUX.1-Kontext-dev lags in instruction following, whereas Nano Banana shows no single dominant weakness.
@@ 1 Introduction
What truly defines the success of an image editor? At its core, editing requires making targeted, instruction-driven changes while preserving contextual consistency and perceptual realism—often across multiple refinement turns. Yet current evaluation practice struggles to capture this multifaceted objective.
@@
When ground-truth edited images are available, a common strategy is to compare model outputs against these references (e.g., MagicBrush Zhang et al. (2023), UltraEdit Zhao et al. (2024), AnyEdit Yu et al. (2025), EmuEdit Sheynin et al. (2024)). Typical metrics include pixel-level distances (e.g., L1/L2) and semantic similarities (e.g., DINO Caron et al. (2021) and CLIP Radford et al. (2021)). While informative, such metrics suffer from two structural issues: (i) the space of acceptable edits is inherently large, whereas a single reference provides only one realization; and (ii) references are frequently synthesized by existing editing models (e.g., Prompt-to-Prompt Hertz et al. (2023), SDXL Podell et al. (2024), DALLE-2 Ramesh et al. (2022)), thereby importing their biases and limitations into the evaluation itself. Consequently, high reference similarity does not necessarily imply faithful instruction following, preservation of irrelevant content, or aesthetically plausible outcomes.
@@
A complementary line of work employs zero-shot VLMs as interpretable evaluators (e.g., VIEScore Ku et al. (2023), GEdit-Bench Liu et al. (2025), I2EBench Ma et al. (2024), HQ-Edit Hui et al. (2024), Complex-Edit Yang et al. (2025), and ImgEdit Ye et al. (2025)) and queries VLMs about specific aspects of an edit. While VLMs offer holistic, language-mediated judgments, they remain insufficient for precise editing assessment for several reasons. First, for instruction-following evaluation, they are notoriously poor at spatial reasoning Zhang et al. (2025b); Chen et al. (2024); Chang et al. (2025) and are prone to object hallucinations in existence, category, attributes, and relations Bai et al. (2024). These issues together undermine their ability to assess common object-related edit instructions. Second, they have limited sensitivity to pixel-level changes and frequently miss subtle, localized modifications Vo et al. (2025) (e.g., fine structures, small attribute shifts, etc.), which are crucial for evaluating content consistency. Third, since they are predominantly pretrained on natural images rather than synthetic generations, their priors are miscalibrated for artifacts and aesthetics, leading to failures in detecting common generative defects (e.g., extra fingers) and in modeling perceptual “naturalness” Liang et al. (2024); Xu et al. (2023); Ma et al. (2025), which humans are sensitive to. Consequently, VLM-only scoring lacks the precision and reliability required for fine-grained evaluation across instruction following, content consistency, and visual quality. However, we find recently open-source state-of-the-art editing models (e.g., Qwen-Image-Edit Wu et al. (2025a), Step1X-Edit Liu et al. (2025), BagelDeng et al. (2025)) solely rely on VLMs for evaluation.
@@
Figure 2: Framework of EdiVal-Agent. It first decomposes images into semantically meaningful objects, such as metal yellow sign and metal brown pole, and identifies their contextual relationships, e.g., they are both in foreground. It then generates diverse and proper editing scenarios at scale which are based on the initial analysis, e.g., Change the color of metal brown pole to gray. Finally, it systematically evaluates editing model outputs from multiple axes with our proposed metrics: EdiVal-IF, EdiVal-$C$, and EdiVal-VQ. Our agentic pipeline is agnostic to the expert tools used and can be readily enhanced with more advanced tools in the future.
@@
To address these challenges, we introduce EdiVal-Agent: an automated and fine-grained evaluation agent for multi-turn instruction-based image editing from an object-centric perspective, designed to assess not only standard single-turn but also multi-turn instruction-based editing with precision. As shown in Fig. 2, EdiVal-Agent first decomposes it into semantically meaningful objects, then synthesizes diverse, context-aware editing instructions while dynamically updating object pools across turns. These two stages enable two novel object-centric metrics tailored for multi-turn evaluation and one global metric of visual quality: 1) EdiVal-IF, which measures instruction following by combining open-vocabulary object detectors for symbolic checks with VLMs for semantic verification on detector-guided crops; 2) EdiVal-CC, which evaluates content consistency by calculating semantic similarity of unchanged objects and background using the evolving object pools; and 3) EdiVal-VQ, which quantifies changes in overall visual quality with human preference models. We show that EdiVal-IF yields stronger agreement with human judgments in instruction-following evaluation compared to thresholded CLIP directional (CLIP_dir) scores Gal et al. (2022) and using VLMs alone, as evidenced in Sec. 2.5.
@@
Instantiating the agentic pipeline, we curate a new multi-turn image editing benchmark, EdiVal-Bench, featuring 9 instruction types and 13 existing editing models—spanning in-context, flow-matching, and diffusion paradigms, across both closed- and open-source systems—conduct fine-grained analyses, and draw actionable insights. Empirically, as demonstrated in Fig. 1 and Tab. 3, GPT-Image-1 excels at instruction following yet ranks near the bottom in content consistency, whereas Seedream 4.0 and Nano Banana performs strongly on both axes. Besides, open-sourced models like Qwen-Image-Edit significantly degrade in instruction following and visual quality when editing turns increase, while FLUX.1-Kontext-max and FLUX.1-Kontext-dev lags in instruction following. We further contrasts multi-turn editing with single-shot complex prompts Yang et al. (2025), highlighting complementary strengths and failure modes. We hope that our agent pipeline, benchmark, and analyses accelerate the transition of multi-turn editing toward practical applications.
Table 1: Key attributes of open-source edit benchmarks. Note that ImgEdit Ye et al. (2025) does not include multi-turn editing experiments in the paper.
Benchmark # Size Object-centric Automated Multi-Turn Free from Ref. Images Tools used
EmuEditSheynin et al. (2024) 3,055 L1, CLIP, DINO
MagicBrushZhang et al. (2023) 1,053 L1, L2, DINO, CLIP
AnyEditYu et al. (2025) 1,250 L1, CLIP, DINO
I2EBenchMa et al. (2024) 2,240 VLM
GEdit-BenchLiu et al. (2025) 606 VLM
HQ-EditHui et al. (2024) 1,651 VLM
ImgEdit-BenchYe et al. (2025) 811 VLM
Kontext-BenchLabs et al. (2025) 1,026 Human Annotation
EdiVal-Bench (ours) 1,716 Detector, VLM, L1, DINO, HPS
@@
Key contributions. 1) Agent: EdiVal-Agent is a fully automated evaluator that performs object-centric decomposition, generates diverse multi-turn editing instructions, and measures overall editing quality using two object-centric metrics (EdiVal-IF and EdiVal-CC) plus EdiVal-VQ for visual quality. 2) Benchmark: using EdiVal-Agent, we construct EdiVal-Bench with 1,716 instructions across nine types and three turns on 572 real-world images, with comparisons to prior benchmarks in Tab. 1. 3) Human agreement: EdiVal-IF attains 81.3% agreement with human ratings for instruction following, outperforming zero-shot VLMs and CLIP-based baselines. 4) Evaluation: we assess 13 editors (diffusion, flow-matching, and close source) along instruction following, content consistency, and visual quality. 5) Insights: overall ranking—Seedream 4.0 >> Nano Banana >> FLUX.1-Kontext-max >> GPT-Image-1; the strongest open-source editor, Qwen-Image-Edit, exhibits exposure bias under multi-turn editing. 6) Artifacts & settings: we reveal luminance drift across turns, and contrast multi-turn against complex single-shot editing to delineate strengths and weaknesses across model families.
2  EdiVal-Agent
@@ 2.1 Overview
As illustrated in Fig. 2. The pipeline comprises three stages: (1) Decomposition uses a VLM (e.g., GPT-4o; other VLMs are viable alternatives) to extract structured, object-level descriptions—objects, attributes, and relations—enabling symbolic reasoning; (2) Instruction Generation produces multi-turn, diverse, compositional prompts by maintaining an explicit object pool and sampling from nine instruction types spanning subject-, attribute-, relational-, text-, count-, and global-level edits; (3) Evaluation scores edited images with EdiVal-IF, Edival-CC, and EdiVal-VQ.
@@ 2.4 Step 3: Evaluation
The first two stages enable two novel object-centric metrics for multi-turn editing evaluation for instruction following and content consistency, respectively, and one global metric for visual quality:
EdiVal-IF
To evaluate instruction following, we introduce EdiVal-IF, which assesses multi-turn edits by comparing the image from the previous turn, $I^{t-1}$, to the current image, $I^{t}$. For a given instruction $P^{t}$ at turn $t$, the score is determined differently for symbolically and semantically verifiable tasks. Symbolically verifiable types ($T_{\text{sym}}$)—such as subject_add, subject_remove, subject_replace, position_change, and count_change—are evaluated using an open-vocabulary object detector $\mathcal{M}_{\text{detect}}$ Liu et al. (2024b). The detector’s outputs, including bounding boxes and confidence, are assessed against geometric and logical criteria $\mathcal{F}_{\text{sym}}$ derived from the instruction. For example, for a position_change instruction “Move [A] to the left of [B]”, $\mathcal{F}_{\text{sym}}$ verifies that the x-coordinate of A’s bounding box $\mathcal{B}$ center is less than that of B in $I^{t}$, i.e., $\text{center}_{x}(\mathcal{B}_{A}^{t})<\text{center}_{x}(\mathcal{B}_{B}^{t})$. In this case,
$\text{EdiVal-}{\text{IF}}(I^{t},I^{t-1},P^{t}\in T_{\text{sym}})=\mathcal{F}_{\text{sym}}(\mathcal{M}_{\text{detect}}(I^{t-1},I^{t}|P^{t})).$ (1)
Semantically verifiable types ($T_{\text{sem}}$)—color_alter, material_alter, text_change, and background_change—are evaluated with a VLM $\mathcal{M}_{\text{VLM}}$ Yang et al. (2024). To focus the evaluation, the VLM is applied to detector-guided object crops ($I_{\text{o}}$) using instruction-specific templates.
$\text{EdiVal-}{\text{IF}}(I^{t},I^{t-1},P^{t}\in T_{\text{sem}})=\mathcal{M}_{\text{VLM}}(I^{t-1}_{\text{o}},I^{t}_{\text{o}}|P^{t})=\mathcal{M}_{\text{VLM}}(\mathcal{M}_{\text{detect}}(I^{t-1},I^{t}|P^{t})).$ (2)
We show that EdiVal-IF achieves superior human agreement (Sec. 2.5). The multi-turn editing success rate is defined as the logical AND of the EdiVal-IF scores across all edits along the chain, whereas the marginal task rate at turn $t$ is defined according to the formulas 1 and 2 provided above.
(a) Base image
(b) GPT-Image-1
(c) FLUX.1-max
Figure 3: Beautification vs. preservation under the prompt: “Change the background to a library.” GPT-Image-1 tends to increase HPSv3 via beautification, while FLUX.1-Kontext-max emphasizes fidelity to the input.
EdiVal-$C$
To assess content consistency, EdiVal-CC measures the preservation of non-target content between the base image $I^{0}$ and the current image $I^{t}$. Given editing instructions $P^{1:t}$ from turn 1 to turn $t$, the object pools $\mathcal{P}^{\text{unch}}_{t}$ and $\mathcal{P}^{\text{all}}_{t}$ are dynamically updated. Let $\Omega$ denote the entire image area. Using object bounding boxes from the base image ($\mathcal{B}^{0}_{o}$) and the current image ($\mathcal{B}^{t}_{o}$), extracted by the detector $\mathcal{M}_{\text{detect}}$, the background region is defined as $\Omega_{\text{bg}}^{t}=\Omega-\bigcup_{o\in\mathcal{P}^{\text{all}}_{t}}(\mathcal{B}^{0}_{o}\cup\mathcal{B}^{t}_{o})$, i.e., the region obtained by excluding all objects that have appeared. Background consistency is then computed as $s^{t}_{\text{bg}}=\phi(I^{0}_{\text{bg}},I^{t}_{\text{bg}})$, where $I^{t}_{\text{bg}}=\Omega_{\text{bg}}^{t}\circ I^{t}$ denotes the background of the image, and $\phi$ is a similarity function such as $L_{1}$ distance or DINO-based similarity. For unchanged objects, we compute the per-object consistency $s^{t}_{o}=\phi(I^{0}_{o},I^{t}_{o})$ for each $o\in\mathcal{P}^{\text{unch}}_{t}$, and then average them. Formally, the final EdiVal-CC score emphasizes semantic preservation by averaging the feature-level similarities of the background and unchanged objects (see Appendix. I.3 for details):
$\text{EdiVal-CC}(I^{t},I^{0},P^{1:t})=\tfrac{1}{2}\left(s_{\text{bg}}^{t}+\tfrac{1}{|\mathcal{P}^{\text{unch}}_{t}|}\sum_{o\in\mathcal{P}^{\text{unch}}_{t}}s^{t}_{o}\right).$ (3)
EdiVal-CC aligns with the intuitive notion of consistency, providing a precise measurement.
EdiVal-VQ.
Zero-shot VLMs are not trained for reliable assessment of image quality—particularly artifacts and aesthetics—and we find they are imprecise as scoring functions (see Appx. H). Consequently, we adopt Human Preference Score v3 (HPSv3) Ma et al. (2025) as our visual-quality (VQ) evaluator. In practice, applying preference models to unedited, real photographs often yields relatively low aesthetic scores. We also observe divergent behaviors across editors (See Fig. 3): some (e.g., GPT-Image-1) tend to beautify images and increase HPSv3, whereas others (e.g., FLUX.1-Kontext-max) preserve the original appearance with minimal aesthetic change. Because aesthetic preference is inherently user- and task-dependent, and beautification can trade off with content consistency (already incorporated into our overall score), we report EdiVal-VQ separately and do not fold it into the aggregate metric.
EdiVal-O.
We aggregate Instruction Following (EdiVal-IF) and Content Consistency (EdiVal-CC) into a single overall score. Since both metrics are unit-free and normalized to $[0,1]$ but capture complementary aspects, we follow prior work and use the geometric mean to balance them and penalize imbalance (Liu et al., 2025; Ku et al., 2023). Formally, $\text{EdiVal-O}\;=\;\sqrt{\text{EdiVal-IF}\;\cdot\;\text{EdiVal-CC}}\,.$
@@
We omit style_change from our taxonomy because style categories are inherently ill-defined, which makes instruction-following (EdiVal-IF) difficult to evaluate reliably. Extending EdiVal-Agent with style-aware recognition is promising future work. After language-based extraction, we validate objects using Grounding-DINO Liu et al. (2024a) and prune low-confidence or ambiguous detections. This stabilizes the object pool and reduces error propagation during instruction generation and IF evaluation. By default, we employ Grounding-DINO as the open-vocabulary detector, Qwen2.5-VL as the VLM, and DINOv3 Siméoni et al. (2025) as the image feature extractor due to their state-of-the-art performance and open-source availability, which facilitates community use. The agentic pipeline is tool-agnostic and can be readily strengthened by substituting more advanced modules in the future.
2.5 Measuring Human Agreement
Figure 4: Results of human agreement. Dashed lines represent the average accuracy of each method. EdiVal-IF achieves 81.3% human agreement accuracy, significantly outperforming the VLM (Qwen2.5-VL) at 75.2% and thresholded CLIP_dir at 65.4%. Note that the CLIP_dir threshold is tuned separately for each task.